3) (a) Suppose, I have normalized the wave function at some point of time. The wave function evolves with time according to time dependent Schrodinger equation. How do


I know that the wave function remains normalized after some time?


[Hint: Show that (x, t)|²dx = 0]


(b) Show that


d(p)


dt


=


dx


i.e. expectation values follow Newton's law.


1
Expert's answer
2021-05-17T13:32:08-0400

Suppose we have a normalized wave function at time t = 0

"\dfrac{d}{dt}\int_{-\infin}^{\infin}|\Psi(x,t)|^2dx=\int_{-\infin}^{\infin}\dfrac{\partial}{\partial t}|\Psi(x,t)|^2dx"


Writing Schrodinger equation as

"\dfrac{d\Psi}{dt}=\dfrac{i\hbar}{2m}\dfrac{\partial^2\Psi}{\partial x^2}-\dfrac{i}{\hbar}V\Psi"


Similarly,

"\dfrac{\partial\Psi^*}{\partial t}=-\dfrac{i\hbar}{2m}\dfrac{\partial^2\Psi}{\partial x^2}+\dfrac{i}{\hbar}V\Psi"


Substituting in the first equation and rearranging, we get,

"\Rightarrow \dfrac{\partial|\Psi|^2}{\partial t}=\dfrac{i\hbar}{2m}\bigg(\Psi^*\dfrac{\partial^2\Psi}{\partial x^2}-\Psi\dfrac{\partial^2\Psi^*}{\partial x^2}\bigg)"

"\Rightarrow \dfrac{\partial|\Psi|^2}{\partial t}=\dfrac{\partial}{\partial x}\bigg[\dfrac{i\hbar}{2m}\bigg(\Psi^*\dfrac{\partial^2\Psi}{\partial x^2}-\Psi\dfrac{\partial^2\Psi^*}{\partial x^2}\bigg)\bigg]"

"\Rightarrow \dfrac{d}{dt}\int_{-\infin}^{\infin}|\Psi(x,t)|^2dx=\dfrac{i\hbar}{2m}\bigg(\Psi^*\dfrac{\partial^2\Psi}{\partial x^2}-\Psi\dfrac{\partial^2\Psi^*}{\partial x^2}\bigg)_{-\infin}^{\infin}"


Since, "\Psi(x,t)\longrightarrow0" as "x\longrightarrow\infin" for "\Psi(x,t)" to be non normalizable,

it follows

"\dfrac{d}{dt}\int_{-\infin}^{\infin}|\Psi(x,t)|^2dx=0"